Jane has a dress shop. She has space for 200 dresses. She sells only red and blue dresses. After reading several fashion magazines she decides to stock her shelves with 34 more blue dresses than red dresses. How many of each should she stock on her shelves?
Answer by ankor@dixie-net.com(22740)

Jane has a dress shop. She has space for 200 dresses. She sells only red and blue dresses. After reading several fashion magazines she decides to stock her shelves with 34 more blue dresses than red dresses. How many of each should she stock on her shelves?
:
Let x = number of red ones
:
It says,"34 more blue than red", therefore
(x+34) = number of blue ones
:
Red + Blue = 200
x + (x+34) = 200
x + x = 200 - 34
2x = 166
x = 166/2
x = 83 Red
and
83 + 34 = 117 Blue
:
:
Check: 83 + 117 = 200